# Thinking Racial Bias in Fair Forgery Detection: Models, Datasets and Evaluations

### v1

@@ 5.2 Benchmarking Fairness of Face Forgery Detectors
Current face forgery detectors exist racial bias variation. Comparing the least fair detector UCF with the most fair detector SPSL, the former’s URDPD is 4.19 times that of the latter, URDEOdds is 3.39 times, URDEO 4.53 times, and URSTD is 3.85 times, showing significant gap in racial bias. Other detectors also exhibit varying degrees of racial bias. Furthermore, we observe that three frequency-based detectors, SPSL, F3Net, and SRM, consistently demonstrate the smallest racial bias across all fairness metrics. This indicates that although current detectors are not explicitly designed for racial fairness, detectors based on different mechanisms still possess varying degrees of racial fairness. We conduct an in-depth investigation into this in Sec 5.3.
5.3 Analyses and Discussions
5.3.1 Balanced Training Dataset.
Data imbalance across races is a prevalent source of bias. To assess the impact of imbalanced training sets on racial bias, we create a dataset that balances across different racial groups. The dataset is obtained by sampling from the FF++. Because FF++ contains a very few number of African and Indian subjects, we can only select 26 subjects from each race. This balanced dataset is employed as the training set. We adopt the same data preprocessing methods and training parameters as the racially imbalanced training set. Notably, due to the reduced dataset size, the number of epochs is doubled to $20$.
Metrics Xception F3Net RECCE UCF
Unbalanced Balanced Unbalanced Balanced Unbalanced Balanced Unbalanced Balanced
Naive Metric DPD 0.1538 0.0852 0.0764 0.0817 0.1317 0.0836 0.1782 0.0908
DEOdds 0.1672 0.0837 0.0877 0.0630 0.1379 0.0663 0.1655 0.0674
DEO 0.2095 0.1076 0.1030 0.1027 0.1777 0.1027 0.2334 0.1104
STD 0.0563 0.0334 0.0278 0.0329 0.0473 0.0314 0.0635 0.0355
Approach Averaged Metric AADPD 0.1957 0.1030 0.1102 0.0996 0.1644 0.1044 0.2107 0.0961
AADEOdds 0.1674 0.0857 0.0951 0.0691 0.1378 0.0746 0.1655 0.0674
AADEO 0.2099 0.1116 0.1178 0.1149 0.1777 0.1193 0.2334 0.1104
AASTD 0.0721 0.0412 0.0425 0.0393 0.0603 0.0405 0.0773 0.0374
Utility Regularized Metric URDPD 0.1314 0.0741 0.0769 0.0770 0.1158 0.0733 0.1433 0.0723
URDEOdds 0.1069 0.0573 0.0624 0.0511 0.0908 0.0505 0.1069 0.0487
URDEO 0.1437 0.0824 0.0842 0.0900 0.1283 0.0847 0.1614 0.0842
URSTD 0.0482 0.0297 0.0295 0.0302 0.0423 0.0284 0.0523 0.0281
Utility Accuracy 0.5552 0.3984 0.5266 0.3308 0.5011 0.4276 0.5557 0.3682
Table 4: Evaluations on the unbalanced and balanced training dataset.
Tab 4 presents the results. For convenience, we put the results of models trained on unbalanced and balanced datasets together. For both Naive Metric and Approach Averaged Metric, we observe that models trained on balanced datasets generally exhibit lower racial bias. Despite the significant improvement brought by a balanced dataset, these detectors still demonstrate a relatively high racial bias. It is crucial to note that, particularly for F3Net, not all metrics demonstrate improvement. These results indicate that training models on a balanced dataset cannot completely address the issue of racial bias, because there are other factors contributing to racial bias. For the Utility Regularized Metric, we note that the enhancement achieved through a balanced training dataset is less significant or may even diminish compared to the Naive Metric and Approach Averaged Metric. This is because this set of metrics is influenced by performance, and the reduced data volume results in a notable model performance drop. This highlights the advantage of the Utility Regularized Metric, i.e., it can reflect the model’s performance.
Based on the aforementioned observations, we argue that utilizing a race-balanced training dataset may not be a recommended way. The fact that deliberately collecting such balanced data in real-world scenarios usually implies discarding a wealth of available imbalanced datasets. Existing datasets are also seldom racially balanced. Furthermore, employing a racially balanced training set does not guarantee effective mitigation of racial bias. In conclusion, training with race-balanced datasets poses challenges, considering the scarcity of such datasets in real-world scenarios and the limited effectiveness in addressing racial bias. Therefore, it is preferable to utilize some other fair learning methods that exhibit better trade-offs.
Figure 4: Probability score distribution of each race.
Xception Utility(Accuracy) ↑ Fairness ↓
Caucasian Asian African Indian STD AccGap
Best Threshold 0.6530 0.7300 0.5380 0.6660 - -
Real Face 0.7960/0.8964 0.7201/0.8813 0.8449/0.8767 0.7715/0.8888 0.0450/0.0075 0.1248/0.0197
FaceSwap 0.8910/0.8265 0.8949/0.7883 0.7668/0.7373 0.8968/0.8264 0.0552/0.0366 0.1300/0.0892
(a) Performance and fairness with 0.5 and optimal thresholds as threshold respectively. Left of ’/’ 0.5. Right of ’/’ is optimal threshold.
Threshold AADPD AADEOdds AADEO AASTD URDPD URDEOdds URDEO URSTD
0.5 0.1274 0.1274 0.1300 0.0501 0.1551 0.1551 0.1507 0.0607
BEST [36] 0.0544 0.0544 0.0892 0.0220 0.0301 0.0301 0.0497 0.0122
(b) Fairness metric results at different thresholds.
Table 5: Threshold Analysis: performance and fairness at different thresholds.
5.3.2 Setting Different Threshold for Each Race.
The study in [36] suggests that in face recognition, the confidence score distributions vary among different race groups. So, they propose setting different thresholds for different races can enhance fairness as well as improve overall performance. We utilize only the real face and FaceSwap portions of FairFD. Firstly, we plot the probability score distribution for each race in Fig 4, revealing distinct differences in the confidence score distribution across different races. The Asian subset shows a higher frequency of high confidence scores, making it more prone to be classified as a fake face. Therefore, a larger threshold can be set for the Asian subset. On the other hand, the African subset exhibits a lower frequency of high probability scores, making it less likely to be classified as a fake face, allowing for a smaller threshold. Next, we set the optimal thresholds, which are values that maximize the overall performance for each racial subset. Tab 5(a) presents the optimal threshold values, as well as the performance and fairness scores when using threshold 0.5 and the optimal thresholds. We observe that, under the condition of maximizing overall performance, there is a certain degree of reduction in racial bias. Tab 5(b) demonstrates racial bias using a fairness metric, showing that indeed, setting different thresholds can reduce racial bias. This conclusion aligns with the findings in [36] consistently.
5.3.3 Analysis of Frequency-Based Detector.
Fairness Metric Xception F3Net RECCE UCF
RGB Grayscale RGB Grayscale RGB Grayscale RGB Grayscale
Naive Metric DPD 0.1538 0.1309 0.0764 0.0780 0.1317 0.0811 0.1782 0.1324
DEOdds 0.1672 0.1439 0.0877 0.0976 0.1379 0.0840 0.1655 0.1512
DEO 0.2095 0.1789 0.1030 0.1035 0.1777 0.1076 0.2334 0.1828
STD 0.0563 0.0470 0.0278 0.0283 0.0473 0.0295 0.0635 0.0475
Approach Averaged Metric AADPD 0.1957 0.1673 0.1102 0.1147 0.1644 0.1097 0.2107 0.1722
AADEOdds 0.1674 0.1439 0.0951 0.1055 0.1378 0.0900 0.1655 0.1512
AADEO 0.2099 0.1789 0.1178 0.1193 0.1777 0.1195 0.2334 0.1828
AASTD 0.0721 0.0614 0.0425 0.0441 0.0603 0.0410 0.0773 0.0658
Utility Regularized Metric URDPD 0.1314 0.1144 0.0769 0.0802 0.1158 0.0785 0.1433 0.1209
URDEOdds 0.1069 0.0935 0.0624 0.0687 0.0908 0.0603 0.1069 0.0986
URDEO 0.1437 0.1249 0.0842 0.0859 0.1283 0.0876 0.1614 0.1321
URSTD 0.0482 0.0419 0.0295 0.0309 0.0423 0.0294 0.0523 0.0461
Table 6: Comparison of fairness on RGB and Grayscale images.
In our previous findings, we observe that frequency-based methods exhibit lower racial bias compared to spatial-based methods. In this section, we provide a preliminary discussion. Spatial-based methods focus on learning forgery clues in the spatial domain, mainly including color mismatch, textures, shapes, and blending boundaries [42]. On the other hand, frequency-based methods concentrate on learning forgery clues in the frequency domain, especially targeting high-frequency information related to blending boundaries, edges, and textures [12]. Frequency-based methods capture less color information, which is also crucial in distinguishing between different racial groups. Consequently, we make the assumption that frequency-based methods’ racial biases are smaller due to these detectors learning less color information. Therefore, we convert the RGB images of our FairFD dataset into grayscale to eliminate color information but retain frequency domain information, and test multiple detectors. The results in Tab 6 demonstrate that the racial bias of spatial-based detectors decreases, while the racial bias of frequency-based methods does not decrease and the changes are small. Thus, color indeed appears to be a contributing factor that leads spatial-based methods to have higher racial bias compared to frequency-based methods. There is still much exploration to be done in the comparative study of frequency and spatial domains, which can significantly contribute to the development of methods for mitigating racial bias. We leave this avenue of research for future investigation.
6 Conclusion
This paper explores how to evaluate the racial fairness of face forgery detectors. We identify numerous disadvantages in existing datasets and fairness metrics, then propose a novel dataset FairFD and two sets of fairness metrics to address these issues. Emphatically, we evaluate the fairness of multiple existing face forgery detectors. The results indicate varying levels of racial bias in current detectors. Further analyses reveal some insights into the emergence of racial bias. In summary, addressing fairness is a crucial progress for the advancement of face forgery detection. We propose several unsolved problems here and hope that our work will help to make progress towards solving them. 1). Although we provide an image-level evaluation dataset, video-level evaluation datasets would still be important. This is because some forgery methods are only applicable at the video level [43, 44, 45], and there are already some detectors based on videos [46, 30] that also suffer from racial bias issues. 2). Although we propose two sets of novel fairness metrics to address the limitations of existing metrics, developing a unified and theoretic metric is still an interesting research direction. 3). Developing fair face forgery detectors that have low racial bias is urgently needed.
Appendix A Appendix
A.1 Classification of Forgery Approaches
There are several approaches for creating fake faces. Here, we provide a brief overview of the classification of forgery methods:
1. Identity-replaced Forgery Approach refers to substituting the original identity in an image, e.g., FaceSwap. FaceSwap [21, 22, 47, 48] involves replacing the face of a person in a video or image with another person’s face. The method usually uses deep learning algorithms to detect and extract the faces of the two people and then swaps them.
2. Identity-remained Forgery Approach retains the original identity and alters other facial attributes, e.g., Attribute Manipulation and Expression Reenactment. Attribute Manipulation [49, 50, 25, 51, 52] involves manipulating the attributes of a face, such as skin color, gender, and eye shape, while preserving the identity of the person. The method usually adopts a GAN. Expression Reenactment [53, 45, 54, 23, 24] involves transferring the facial expressions of one person to another person’s face. The method usually detect and track the facial landmarks of the two people and then transfers the expression from the source face to the target face.
A.2 Existing Fairness Metrics
Due to the complexity of fairness evaluation, numerous fairness metrics are proposed from various perspectives to cater to different scenarios. Therefore, we need to leverage multiple fairness metrics to evaluate the fairness in face forgery detection comprehensively. Here, we present the four most commonly used fairness metrics and outline their respective applicable scenarios.
Demographic Parity Difference (DPD): In the context of face forgery detection, where label 1 represents fake face, DPD reflects the model’s inclination to categorize faces of a specific race as fake. When people perceive the classification of faces as fake as a form of discrimination, we can leverage DPD to assess the extent of bias in the model.
Difference in Equalized Odds (DEOdds): DPD may fail in certain situations [55]. Considering a scenario where a face forgery detector classifies faces of African and Caucasian individuals as fake at a similar rate, but the model makes different types of errors for the two groups. Specifically, for African faces, the false positive rate is significantly higher than for Caucasian faces, while for Caucasian individuals, the false negative rate is higher. In such a case, we can use DEOdds.
$DEOdds=\frac{1}{2}\underset{y=\left\{0,1\right\}}{\Sigma}\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}|Y=y,S=s\right)-P\left(\hat{Y}|Y=y,S=s^{\prime}\right)\right|.$ (5)
Equal Opportunity (DEO): DEO has more relaxed conditions compared to DEOdds. When assessing the fairness between group A and group B, only the images of faces that are inherently fake need to be considered. DEO solely focuses on true positive rates and does not capture the overall classification differences.
$DEO=\max_{s,s^{\prime}\in\mathbb{S},s\neq s^{\prime}}\left|P\left(\hat{Y}|Y=1,S=s\right)-P\left(\hat{Y}|Y=1,S=s^{\prime}\right)\right|.$ (6)
Standard Deviation (STD): STD differs from metrics mentioned above. STD considers the overall variability rather than just the differences between the best and worst-performing ethnicities. We use accuracy as the performance metric.
$STD=std\left(\left\{acc\left(S=s\right)_{s\in\mathbb{S}}\right\}\right),$ (7)
where $std$ is a function that calculates the standard deviation of given list. $acc$ calculates the accuracy of a given race.
A.3 Face Forgery Detection Datasets
We summarize the information of existing datasets in Tab 2. Below we give simple description of some widely used face forgery datasets.
FaceForensics++(FF++) [1] is a forensics dataset that consists of 1000 original video sequences downloaded from the Internet(,i.e.,YouTube). The videos are manipulated with four automated face manipulation methods. For face swap, they use Deepfakes [56] and FaceSwap [21]. For expression reenactment, they use Face2Face [45] and NeuralTextures [57].
UADFV [26] contains videos of varying classes, with each video being classified as either real or fake. The dataset is relatively small, with only 98 videos, but it has been found to be convenient in terms of how the data is formatted.
CelebDF-v2 [27] is a comprehensive collection for deepfake forensics, comprising 590 real videos and 5,639 DeepFake videos featuring celebrities with high-quality. These videos were generated through an enhanced synthesis process, ensuring superior visual quality and better representing the DeepFake content prevalent on the internet.
Deepfake Detection Challenge(DFDC) [28] is created by Facebook in partnership with other industry leaders and academic experts. The dataset consists of 128,154 videos featuring 960 paid actors. The dataset was used in a Kaggle competition to create new and better models to detect manipulated media.
DeeperForensics-1.0(DF-1.0) [29] contains 60,000 videos and 17.6 million frames with 100 consented actors. The actors in DF-1.0 have four skin tones: white, black, yellow, brown, with roughly balanced ratio. Different from other datasets, DF-1.0 attach importance to high-quality and diversity of source face videos. Each actor has various poses, expressions, and illuminations. And DF-1.0 has more real videos than fake videos with a ratio of 5:1.
ForgeryNet [30] is a very large dataset for real-world face forgery detection, with 2.9 million images, 221,247 videos and 15 forgery approaches. And a pipeline of conducting various face forgery approaches are proposed.
A.4 Face Forgery Detection Algorithms Categories
Current face forgery detectors can be roughly divided into two categories: Spatial-based, and Frequency-based. Spatial-based method [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] is based on the spatial domain features(,i.e., forgery clues) of the fake face. Such as RECCE [10] utilizes an encoder to reconstruct real face images, thereby exploring the differences between real and fake images in the spatial domain. The differences in the reconstruction of the two are then used as guidance to train a classifier for detecting deepfakes. Frequency-based method [12, 13, 14] identifies distinctions between real and fake faces in frequency domain, which are used to detect whether the face is forged. Such as F3Net [12] utilizes frequency-aware decomposed image components and local frequency statistics to explore forgery patterns, enabling effective forgery detection.
A.5 Fairness in Face Recognition
An increasing number of researchers are are now paying attention to societal issues of artificial intelligence, including adversarial example [58, 59, 60], privacy protection [61, 62, 63], and fairness concerns.
The research on fairness in face forgery detection is similar to the study of fairness in face recognition. RFW [32] is first introduced as a race balanced test dataset for face recognition. They also propose IMAN which uses a deep information maximization adaptation network to align global distribution to decrease race gap at domain-level, and learns the discriminative target representations at cluster level. Another balanced face recognition dataset BFW [36] is proposed to evaluate the fairness of face recognition system both for gender and ethnic groups and this work also shows variations in the optimal scoring threshold for face-pairs across different subgroups.
The issue of racial bias in machine learning has garnered significant attention in multi fields. The research on fairness in face forgery detection is akin to the study of fairness in face recognition. RFW [32] is first introduced as a balanced test dataset for ethnic group. IMAN [32] uses a deep information maximization adaptation network to align global distribution to decrease race gap at domain-level, and learns the discriminative target representations at cluster level. Another balanced faces dataset BFW [36] is proposed to evaluate the fairness of face recognition system both for gender and ethnic groups and this work also shows variations in the optimal scoring threshold for face-pairs across different subgroups. DebFace [37] uses a de-biasing adversarial network to extract disentangled feature representations for both unbiased face recognition and demographics estimation and adopts adversarial learning to minimize correlation among feature factors so as to abate bias influence. [38] involves multiple preprocessing methods to improve the dual-shot face detector, data re-sampling to balance the data distribution, and multiple data enhancement methods to increase accuracy performance and proposes a linear-combination strategy is adopted to benefit from multi-model fusion. Meta Balanced Network [64] uses meta-learning algorithm to learn adaptive margins in large margin loss to mitigate the algorithmic bias in face recognition models. MixFairFace [39] proposes MixFair Adapter to determine and reduce the identity bias of training samples.
A.6 Number of Subjects is Limited
We propose that the limited number of subjects makes it challenging to draw meaningful comparisons across groups.
Caucasian Asian African Indian
TPR 0.7764 0.7365 0.6289 0.6801
TNR 0.9502 0.9652 0.9314 0.9646
Table 7: TPR and TNR of each race.
Here we do a verification experiment to explain and prove it. We utilize the FF++ dataset, retains only the Caucasian, Asian, African, and Indian subsets. The remaining dataset consists of a total of 677 subjects. For each race, 16 subjects are chosen as the test dataset, and the rest serve as the training dataset (9:1 approximately). We train an Xception model on the training dataset for deepfake detection and evaluate its performance on the test dataset. We follow the hyperparameters specified in [41] and train the model for 10 epochs.
Figure 5: TPR and TNR for each subject.
In Fig 5, we show the TPR and TNR on each subject. We can observe significant fluctuations in the model’s performance across each subject. The presence of extreme results also indicates that subjects significantly influence the model’s classification. Therefore, when the number of subjects is insufficient, it becomes challenging to capture the overall characteristics of a group, making it difficult to draw meaningful comparisons across race groups.
A.7 Visualization in Feature Space
The research on fairness in face forgery detection is similar to the study of fairness in face recognition(see Sec A.5 of Supp for more details). As stated in [32], one reason for racial bias is that different subsets’ features are totally separate. So they take the racial bias as a problem of domain gap and propose IMAN(information maximization adaptation network) to decrease this domain gap. We sample 500 samples for each ethnic group and use the well trained Xception model as the detector to obtain features for these samples. Then, we plot the t-SNE dimensionality reduction graphs for the feature spaces of the four ethnic subsets in Fig 7. Considering that different forgery approaches will cause distinct results, we plot for Real Face, FaceSwap and FaceReen respectively. Unlike in [32], we do not observe distinct separation between different subsets at feature level. Next, we use the MMD(Maximum Mean Discrepancy) to mathematically calculate the feature distances between the Caucasian subset and other ethnic groups. The results in Fig 7 demonstrate that although distances show difference, in comparison to the distances in [32], the distances here are all nearly close to zero. These results are because they consider the face recognition task in [32], so the models tends to learn distinctive features for each subject, resulting in significant differences at the feature level. In contrast, for the task of face forgery detection, the detector does not exhibit the same tendency, leading to smaller differences in feature level. Therefore, attempting to enhance the fairness of face forgery detection from a feature-level perspective may not be feasible.
Figure 6: T-SNE visualization.
Figure 7: MMD.

### v3

@@ Introduction
Benchmark Results. The benchmark results (shown in Table 2) present a comprehensive evaluation of 12 face forgery detectors using various fairness metrics. We highlight the four fairest detectors using different colors. Based on the results, we draw the following significant observations: (1) Current face forgery detectors all exhibit a high degree of racial bias. The DPD metric of SPSL can achieve 0.0203 with the smallest racial bias. This indicates that the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 2.03%. When separately calculating and averaging for each forgery method, the AADPD metric reaches 5.56%. On the other hand, for the least fair detector UCF, the difference in the probability of classifying faces as fake between the most advantaged and disadvantaged groups is 17.65%. This reminds researchers to address the racial bias in existing face forgery detection models. (2) Current face forgery detectors have racial bias variation. Comparing the least fair detector UCF with the most fair detector SPSL, the former’s URDPD is 4.76 times that of the latter, URDEOdds is 3.58 times, URDEO 4.98 times, and URSTD is 4.48 times, showing significant gap in racial bias. Other detectors also exhibit varying degrees of racial bias. Furthermore, we observe that three frequency-based detectors, SPSL, F3Net, and SRM, consistently demonstrate a smaller racial bias across all fairness metrics. We conduct an in-depth investigation into this in the section "Analyses and Discussions" in Supp.
Detailed Utility Results. To present more detailed results, we present the AUC for each detector, each race, and each forgery method in Figure 3. Results show that different forgery methods exhibit varying levels of utility. This validates the advantage of Utility Regularized Metric.
Detailed Fairness Results. We present the standard deviation (STD) of the ACC for the four races in Figure 4 for each forgery method(including Real Face). Our findings reveal that different forgery methods exhibit varying levels of fairness, and different detectors rank the fairness of these forgery methods differently. This validates the advantage of the proposed Approach Averaged Metric.
Figure 4: Fairness (STD) for different forgery approaches.
Evaluating BPFA
Baseline Algorithm
We select two baseline methods for comparison. The first WEIG uses only the absolute values of the weights as the pruning score, and the second RoBA uses only the reciprocal of the bias of the activations. More details about these baselines are shown in Supp. For all three methods, we prune the parameters with the lowest pruning scores. These pruning baselines can be directly applied in existing SOTA forgery detection models.
Results Analysis
The experimental results under optimal pruning rates for each detector and method are shown in Table 3. It can be found that the proposed method BPFA consistently outperforms all baseline methods, achieving superior fairness without compromising utility. Although WEIG generally preserves good utility and enhances fairness, its improvements in fairness are not as significant as those achieved by BPFA. On the other hand, RoBA exhibits highly unstable performance. It results in improving fairness but at the cost of significantly reduced utility, which can render the model nearly unusable in the forgery detection task. These results prove the superior performance of BPFA in enhancing both utility and fairness for forgery detection. It is encouraging to find that SPSL+BPFA achieves the new state-of-the-art performance. The only parameter in our method is the pruning rate. The ablation study for pruning rate is detailed in three tables in the section "Ablation Study on Pruning Rate" in Supp. Our findings indicate that different detectors and different methods have significantly different optimal pruning rates (which correspond to the least decrease in utility (or even improvement) while achieving the best average value across the 12 fairness metrics.
Analyses and Discussions
Here we conduct deeper analyses and give some insights: (1) We train detectors using balanced training data, finding that while racial bias can be reduced, the cost of collecting balanced data is substantial; (2) We set the optimal classification threshold for each race and then use the resulting accuracy values to calculate fairness. The conclusions show that this method is highly cost-effective and significantly improves both utility and fairness simultaneously; (3) We analyze that frequency-based detectors exhibit superior fairness performance because of not utilizing race-sensitive information, e.g. color. More analyses and details are shown in Supp.
Conclusion
This paper early explores a comprehensive racial bias evaluation benchmark for forgery detection, which provides a newly self-construct dataset, fairness metric and unified protocols. We identify numerous disadvantages in existing datasets and fairness metrics, then propose a novel dataset FairFD dataset and two sets of fairness metrics to address these mentioned issues. Besides, we also propose a novel Bias Pruning with Fair Activations algorithm to improve the fairness performance without an extra training process. Emphatically, we evaluate the fairness of multiple existing face forgery detectors. The results indicate the racial bias in current detectors is generally high and prove the advantages of our proposed BPFA. Further analyses reveal some interesting insights into the emergence of racial bias. We hope the proposed benchmark can inspire more researchers to develop the field. In the future, we will explore a unified fairness metric for diverse biases in more kinds of datasets, and construct the video-level forgery detection datasets for more real applications. See social impact in Supp.
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62306227, Grant 62276198, Grant U22A2035, Grants U22A2096, Grant 62441601 and Grant 62036007; in part by the Fundamental Research Funds for the Central Universities under Grant ZYTS24142, Grant QTZX23083 and Grant QTZX23042; in part by the Key Research and Development Program of Shaanxi (Program No. 2023-YBGY-231); in part by Young Elite Scientists Sponsorship Program by CAST under Grant 2022QNRC001; in part by the Guangxi Natural Science Foundation Program under Grant 2021GXNSFDA075011; in part by the Shaanxi Province Core Technology Research and Development Project under grant 2024QY2-GJHX-11; in part by Open Research Project of Key Laboratory of Artificial Intelligence Ministry of Education under Grant AI202401, in part by the Nanning Scientific Research and Technological Development Project 20231042; in part by the ‘111 Center’ (B16037).